OpenStudy (anonymous):

simplify: 2a+2/a^2-1

OpenStudy (aravindg):

simple take common factors out in numerator

OpenStudy (aravindg):

In denominator use identity \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (gorv):

a^2-1=(a-1)(a+1) 2(a+1)/(a+1)(a-1)=2/(a-1)

OpenStudy (aravindg):

@gorv giving direct worked out answers is against openstudy code of conduct refer http://openstudy.com/code-of-conduct

OpenStudy (aravindg):

Instead help the asker get to the solution being a guide along the process

OpenStudy (anonymous):

okay to the answer would be 2/a-1 right?